OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

Take the first equation an solve for x x=(1-y)/2. Then substitute this x into the second equation and solve for y 2=-2 However, we run into a contradiction, and the system has no solutions.

OpenStudy (anonymous):

2x + y = 1 => 2x = 1 - y 4x + 2y = −2 = 2(2x) + 2y = -2 substituting for 2x this becomes 2(1 - y) + 2y = -2 2 - 2y + 2y = -2 2 = -2 this is a contradiction, so this system of equations does not have a solution